The Gilster Company, a machine tooling firm, has several plants. One plant, located in St. Falls, Minnesota, uses a job order costing system for its batch production processes. The St. Falls plant has two departments through which most jobs pass. Plantwide overhead, which includes the plant manager’s salary, accounting personnel, cafeteria, and human resources, is budgeted at $250,000. During the past year, actual plantwide overhead was $242,000. Each department’s overhead consists primarily of depreciation and other machine-related expenses. Selected budgeted and actual data from the St. Falls plant for the past year are as follows.

Department A Department B
Budgeted department overhead
(excludes plantwide overhead) $ 97,500 $ 540,600
Actual department overhead 120,000 558,600
Expected total activity:
Direct labor hours 56,000 15,000
Machine-hours 13,000 53,000
Actual activity:
Direct labor hours 58,500 14,000
Machine-hours 13,700 55,000

For the coming year, the accountants at St. Falls are in the process of helping the sales force create bids for several jobs. Projected data pertaining only to job no. 110 are as follows.

Direct materials $ 21,500
Direct labor cost:
Department A (2,400 hr) 36,000
Department B (800 hr) 11,600
Machine-hours projected:
Department A 100
Department B 1,200
Units produced 12,000

Required:

a-1. Assume the St. Falls plant uses a single plantwide overhead rate to assign all overhead (plantwide and department) costs to jobs. Find the overhead rate by using expected direct labor hours.

a-2. Determine the projected amount of total manufacturing costs per unit for the units in job no. 110.

Assume the St. Falls plant uses three separate overhead rates to assign overhead costs to jobs.

b-1. Find the plant wide overhead rate by using expected machine hours.

b-2. Find the department overhead rate using expected machine hours for Department A and Department B.

b-2. Calculate the projected manufacturing costs for job 110 using the three separate rates computed in b-1 and b-2.

c-1. The sales policy at St. Falls dictates that job bids be calculated by adding 23 percent to total manufacturing costs. What would be the bid for job no. 110 using the overhead rate from part a?

c-2. The sales policy at St. Falls dictates that job bids be calculated by adding 23 percent to total manufacturing costs. What would be the bid for job no. 110 using the overhead rate from part b?

c-3. Which of the overhead allocation methods would you recommend?
Answer #1
a1) Plantwise overhead rate = Expected Plant overhead / total plant
(250000+97500+540600) / (56000+15000) = 12.50845 per DLH
a2) projected total manuf. Costs of Job 110 =
Amount $
dM 21500
dl 47600
overhead allocation: 3200*12.50845 40027.04
Total cost of Job 110 109127
Units produced 12000
Total Manuf. Cost per unit 9.09
b1) plantwise overhead rate = 250000 / (13000+53000)=3.787879 per MH
b2)
Deptt. A = 97500 / 13000 = 7.5 per MH
Deptt. B = 540600 / 53000 = 10.20 per MH
b3) projected total manuf. Costs of Job 110 =
Amount $
dM 21500
dl 47600
overhead allocation:
Deptt A 100*7.50 750
Deptt B 1200*10.2 12240
Plantwise 1300*3.787879 4924.24
Total cost of Job 110 87014.24
c1)
Bid price:
Total Manufacturing cost 109127
add: Margin 23% 25099.21
Bid price: 134226.21
c2)
Bid price:
Total Manufacturing cost 87014.24
add: Margin 23% 20013.28
Bid price: 107027.52
c3) The allocation through Machine Hour and departmental and  
Plantwise overhead separate allocation is suitable because
actual used department's overhead are allocated to job and plantwise
overhead is allocated according to actual machine hour used. The method
keeps the bid price in competitive mode.